OpenStudy (anonymous):

Solve the equation \[log_4(x)+log_7(x)=1\]

OpenStudy (turingtest):

@purplec16 could you please not post all in caps? It makes it seem as though you are yelling, and is a bit harsh on the eyes. thanks

OpenStudy (turingtest):

also the fastest way to get help is to just type the actual question in the post, so you should probably go ahead and do that

OpenStudy (anonymous):

oh... I wasn't yelling... sorry I forgot above that...

OpenStudy (turingtest):

it's cool, a lot of people do it

OpenStudy (anonymous):

I've had questions up for hours and noone helped me...

OpenStudy (anonymous):

Do you know how to solve \[log_4(x)+log_7(x)=1\]

OpenStudy (turingtest):

I think I have a plan... let me work it out

OpenStudy (anonymous):

Yay! :D

OpenStudy (turingtest):

do you know the change of base log formula?

OpenStudy (anonymous):

I know it... but my lecturer skipped it... and I don't know how to use it...

OpenStudy (turingtest):

@Callisto i have a feeling that you have a better idea than mine, so feel free to put your solution if you have one

OpenStudy (turingtest):

I was doing something like the change of base formula, which can be written as the change of log base a to log base b is\[\log_bx={\log_b x\over \log_b a}\]

OpenStudy (turingtest):

so taking b=e we can make this a natural logarithm\[{\ln x\over\ln4}+{\ln x\over\ln7}=1\]\[\ln7\cdot\ln x+\ln4\cdot\ln x=\ln4\cdot\ln7\]\[\ln x={\ln4\cdot\ln7\over\ln7+\ln4}\]

OpenStudy (anonymous):

how'd you go from the first to second step?

OpenStudy (turingtest):

factor out lnx\[\ln7\cdot\ln x+\ln4\cdot\ln x=\ln4\cdot\ln7\]\[\ln x(\ln7+\ln4)=\ln4\cdot\ln7\]\[\ln x={\ln4\cdot\ln7\over\ln7+\ln4}\]

OpenStudy (turingtest):

oh first to second? the change of base formula I wrote above

OpenStudy (turingtest):

\[\log_bx={\log_b x\over \log_b a}\]

OpenStudy (callisto):

\[{\ln x\over\ln4}+{\ln x\over\ln7}=1\]\[{\ln7\ln x\over\ln4 \ln7}+{\ln4\ln x\over\ln4\ln7}=1\]\[{\ln7\ln x+\ln4\ln x\over\ln4\ln7}=1\]\[\ln7\ln x+\ln4\ln x=\ln4\ln7\]

OpenStudy (turingtest):

oh, that step :P I wasn't sure which one you were talking about I guess

OpenStudy (anonymous):

Yeah and it's the same thing @Callisto... what did you do for your second line and why?

OpenStudy (turingtest):

She just got a common denominator between the two fractions

OpenStudy (callisto):

Addition of fraction... you need to have the same denominator for that, right?

OpenStudy (anonymous):

Yes... so you multiply each... but I am not seeing it when you're skipping steps.. or idk... it's just a little more difficult with logarithm, sigh... I feel so confused and defeated

OpenStudy (callisto):

Multiplying the same thing in numerator and denominator, adding them with the same denominator, multiplying the denominator to both sides.... Which step did I skip? :O Is it clearer to you now??

OpenStudy (anonymous):

actually no... isn't the common denominator ln4ln7...

OpenStudy (callisto):

Which part you don't understand?

OpenStudy (anonymous):

Oh, I just realized the last line is what I was implying to but I didn't understand your method but that's okay... as of now I don't understand how you got the first line and how to solve the last line...

OpenStudy (callisto):

If you don't mind... how do you simplify it? \[\frac{1}{b}+\frac{1}{a}\]

OpenStudy (anonymous):

\[\frac{a+b}{ba}\] I got what you did with the fraction.. but I don't know how to solve the log and I got the first line now too :D

OpenStudy (callisto):

''solve the log''? If you have lnx = 1, how do you solve it?

OpenStudy (anonymous):

e^1=x

OpenStudy (anonymous):

I'm really sorry but I haven't done any of the complicated with lnx*ln7 I don't know what to do with that... does is become ln 7x?

OpenStudy (callisto):

No... lnx * ln7 ≠ ln (7x)

OpenStudy (anonymous):

I don't know how to add two log's...

OpenStudy (callisto):

ln7 + ln 4 = ln (7x4) = ...?

OpenStudy (anonymous):

28

OpenStudy (anonymous):

but... but... :\... why are log's so confusing I can't feel the right side of my head...

OpenStudy (callisto):

I don't know too :|

OpenStudy (anonymous):

So it's lnx^2ln(28)=ln4ln7

OpenStudy (callisto):

No.... you should ''isolate'' x instead.

OpenStudy (anonymous):

by factoring?

OpenStudy (callisto):

No. Like the way you solve lnx =1 Now, it's \(\large lnx = \frac{ln4ln7}{ln4+ln7}\)

OpenStudy (anonymous):

Doesn't that involving factoring the lnx from the left hand side... \[lnx(ln7+ln4)=ln4ln7\] maybe I said it wrong... Idk...

OpenStudy (anonymous):

So this \[x=e^{\frac{ln4ln7}{ln28}}\] In the exponent is...\[\frac{ln4ln7}{ln28}\]

OpenStudy (anonymous):

\[x \approx 2.246908864...\]

OpenStudy (callisto):

Yup!

OpenStudy (anonymous):

Sigh... that was rough... sorry for "yelling" earlier @TuringTest but... I'm preparing for finals and I'm really anxious, stressed, depressed/emotional... and really tired. BUT...Thanks so much @Callisto and @TuringTest I really really really do appreciate your help! Like seriously, you all are awesome!!! I'm so glad I've found you... lol :D.

OpenStudy (callisto):

'' I'm really anxious, stressed, depressed/emotional.'' => That's me!!!! No worries! Everything will be fine!!! Good luck for the finals too!

OpenStudy (anonymous):

Thanks so much! Like I really appreciate you all... because I would have been a nervous wreck if I didn't learn this...and might have forgot how to add/divide/subtract/multiply... lol!!!

OpenStudy (anonymous):

Would you mind helping me very quickly with another one, in another post... it shouldn't be as long because I think I grasped the basic idea...

OpenStudy (callisto):

Just post the question. I'm sure there's always someone to help!

OpenStudy (anonymous):

Okay :D, thanks